OpenStudy (anonymous):

how do i solve this quadratic equation? 9x^2-6x+1=0 i tried using the quadratic formula and i got 6 for my answer but i want to make sure i did it right!

OpenStudy (anonymous):

to make sure you did it right you can check by putting 6 into your equation as x 9(6^2) - 6(6) + 1 = 289 which isn't zero :( show me your workings and ill guide you through it :D

OpenStudy (zzr0ck3r):

9x^2 -6x+1=0 9x^2 - 3x - 3x +1 = 0 3x(3x-1)-1(3x-1) = 0 (3x-1)(3x-1)=0 3x-1 = 0 3x = 1 x=1/3

OpenStudy (anonymous):

\[ok so i did -(-6)\sqrt{-6}^{2}-4(9)(1)\div2(9)\]ok so i did ...-(-6)

OpenStudy (anonymous):

so @zzr0ck3r has shown you the factorizing method of solving it to use the equations you find: a = 9 b = -6 c= 1 then do: \[\frac{-b \pm \sqrt{(b^2 - 4ac)}}{2a}\]

OpenStudy (anonymous):

\[\frac{6 \pm \sqrt{(36 - 36)}}{18}\]

OpenStudy (anonymous):

= 6/18 = 1/3

OpenStudy (zzr0ck3r):

and the best way to do these IMO is completing the square 9x^2-6x+1=0 x^2 -2/3x+1/9 = 0 (x-1/3)^2 = -1/9 +4/36 (x-1/3)^2 = 0 x-1/3 = 0 x = 1/3

OpenStudy (zzr0ck3r):

if you get good at completing the square you can tell the answers to these just by looking at them

OpenStudy (anonymous):

i need to brush up on that, i have got rusty over the holidays

OpenStudy (zzr0ck3r):

first get rid of the third term by moving over to the RHS then devide away any coeff that is on the first term then take half the second term and square it and add it to the RHS then take the (first term -+ the second term cut in half)^2 then solve (the +- depends on the sign of the second term)

OpenStudy (anonymous):

see that just seems like doing the formula to me. in class i tended to just do the formula quickly in my head. but that's just what i found easier.

OpenStudy (zzr0ck3r):

yeah its the same thing in a sense..